OpenStudy (anonymous):

integrate x^2 cos x^3 which method most suitable? by part or substitution??

myininaya (myininaya):

sub! let u=x^3 => du=3x^2 dx

OpenStudy (anonymous):

how about \[\int\limits_{}^{}(x^2)dx/ \sqrt{1- (x-1)^2}\] hint: \[x-1=\sin \theta\]

OpenStudy (mimi_x3):

\[=\int\limits x^{2}cosu*\frac{du}{3x^{2}} \] \[= \frac{1}{3} \int\limits cosu du \] \[= \frac{1}{3} *sinu + c\] \[= \frac{1}{3} *sinx^3 + C\]

OpenStudy (aravindg):

nikvist can u hlp

OpenStudy (anonymous):

i was lurning si and stuff like that